Home Depot entered fiscal 2016 with a total capitalization of $27,252 million. In 2016, debt investors received interest income of $843 million. Net income to shareholders was $6,384 million. (Assume a tax rate of 35%.) Calculate the economic value added assuming its cost of capital is 10%. (Do not round intermediate calculations. Enter your answer in millions rounded to 2 decimal places.)
Answer #1
EVA = NOPAT - (Capital invested*WACC)
NOPAT = 6384+843*(1-35%) = 6931.95
Capital invested * WACC = 27252*10% = 2725.20
EVA = 6931.95-2725.20 = 4206.75